OpenStudy (anonymous):

Prove that x = 2: 3^(x/2) + 1 = 2^x

OpenStudy (anonymous):

3^ (2/2) + 1 = 2^2 this should work think you can try to solve from here?

OpenStudy (anonymous):

Sorry, i should have expressed myself better, i need an expression for the value of x, i mean, isolate x in this expression

OpenStudy (anonymous):

i see, so grouping it should work like 1 =3^(2/2) - 2^2 ( i would still fill in x for 2 because it is already given so what is 3 ^1?

OpenStudy (anonymous):

ok i have this backwards

OpenStudy (anonymous):

Don't worry, i know how to plug in the value. I was just wondering if there is any way to isolate x, but apparently it's not possible. But thanks man!

OpenStudy (anonymous):

I the expression was more complex, maybe the only solution would be a numerical method, such as Newton's method

OpenStudy (anonymous):

i would think that 3 -2 before the exponents would work